# AniFaceDiff: High-Fidelity Face Reenactment via Facial Parametric Conditioned Diffusion Models

### v1

@@ 4.2 Ablation Study
Table 2: Quantitative comparison with or without Facial Shape Alignment and the Expression Adapter. The addition of both components allows for a better trade-off between identity consistency vs expression fidelity. We find that combining both provides the best qualitative and quantitative results.
. Same-identity Reconstruction Cross-identity Reenactment FSA EA CSIM APD AED CSIM APD AED - - 0.75 0.011 0.098 0.52 0.021 0.182 - - - - 0.59 0.022 0.206 - 0.75 0.011 0.091 0.51 0.021 0.158 - - - 0.57 0.022 0.174
Figure 5: Qualitative comparison on cross-identity reenactment with or without Facial Shape Alignment and the Expression Adapter. FSA improves facial generation quality, while EA maintains expression consistency. Ensembling both provides the best overall consistency. These observations are consistent with the quantitative results in Table 2.
We conduct an ablation study on the proposed 1) Facial Shape Alignment (FSA) and 2) Expression Adapter (EA). Facial Shape Alignment is used exclusively for cross-identity reenactment. Our baseline method directly conditions the model with 2D facial snapshots extracted from the driving video, without the facial shape alignment and the expression adapter. As illustrated in Table 2, adding facial shape alignment significantly improves identity preservation (CSIM). Facial shape alignment uses the shape embedding from the source image to render the 2D facial snapshots, thereby enhancing identity preservation. Without facial shape alignment, the output maintains the shape from the driving image rather than the source image, as shown in Fig. 5. The expression adapter significantly enhances expression accuracy (AED) in both same-identity reconstruction and cross-identity reenactment. Methods incorporating the expression adapter can provide more expression-related details, such as wrinkles, and achieve greater accuracy in certain areas, such as the mouth, as depicted in Fig. 5. Combining both FSA and EA can effectively handle the trade-off between identity consistency and expression accuracy.
@@ 5 Ethical Considerations
We are committed to positive virtual face reenactment applications. All photorealistic human portraits presented in this paper are virtual and do not exist in reality. Acknowledging the broad applicability of face reenactment, we are also aware of its potential for malicious misuse, such as in deepfake fraud. To address this concern, we plan to implement stringent access control mechanisms to limit access permissions exclusively for research purposes. Additionally, we evaluate existing SOTA deepfake detectors (see Appendix A.1) and conduct an analysis to identify potential shortcomings. This will serve as a foundation for future research aimed at mitigating weaknesses of existing detectors and providing benchmarking data for future approaches.
Appendix A Appendix / supplemental material
A.1 On deepfake detection and potential for misuse
While any face reenactment technique can also be misused for deepfake generation, it is clearly an important area of active research with many ethical applications in movies, journalism, animation and virtual reality, to name a few. This is especially apparent in emergent capabilities in techniques such as ours, as improved results in the primary domain (real human face animation) can lead to significant improvements in cross-domain or out of domain animation transfer (human to 3D, human to cartoon, cartoon to 3D, cartoon to cartoon etc.) However, to mitigate the deepfake generation potential of our approach, we perform a detailed and comprehensive analysis of state of the art detection techniques on our method, and report their zero shot performance on our data. First, this contributes to the understanding of the generalization capabilities of existing deepfake detectors. Secondly, it highlights potential for improving the training datasets of such detectors, especially with videos based on paradigms using diffusion models as the generative core. This work opens up the possibility of creating a large scale dataset of high fidelity diffusion based face reenactment videos, addressing the issue of lack of such data in the deepfake detection domain.
Deepfake Autoencoder (DFAE, e.g. DeepFaceLab [^2] https://github.com/iperov/DeepFaceLab) and other face swapping methods can produce highly realistic videos that are challenging to detect using existing detection algorithms. This is evident in the results of detection algorithms on the DFDC [64] dataset. Additionally, videos generated by face reenactment methods such as Neural Texture [65] and Face2Face [66] are also difficult to detect using off-the-shelf detectors (details in [67]). It is worth noting that these methods do not offer any alternative means of detecting videos generated by them, which encourages generation of defamatory and harmful deepfakes.
@@ A.1 On deepfake detection and potential for misuse
Table 3: Detection frame-level and video-level AUC from the state-of-the-art deepfake detectors with respective pretrained weights on videos generated by our reenactment method AniFaceDiff.
We emphasize the significance of ensuring that these methods are used ethically and that they can be detected with existing deepfake detectors; this would invalidate videos generated with malicious intent. We evaluated videos generated with our method on pretrained, state-of-the-art deepfake detection methods. Table 3 presents both frame and video-level detection results. Although video-level detection is more common, we include frame-level detection due to the possibility of temporally partial deepfakes [67] i.e. having both real and fake segments in the same video. Our evaluation encompasses multiple types of detectors including naive detectors [68, 70], a frequency-level detector [72] and spatial detectors [69, 71, 73, 74, 75]. We notice that videos generated with AniFaceDiff had high detection rate by the tested detectors. Frame-level detection is more challenging since the detection performance depends on the prediction of each frame, rather than a majority voting across the frames in a video. While the results are promising, there is room for improvement. Further investigation into diffusion-based reenactment methods could further enhance detection rates.
@@
This assessment aims to pave the way for future studies on detection. Additionally, we provide access to researchers from deepfake detection in order to facilitate the generation of higher quality deepfake datasets, particularly those generated using state of the art face reenactment diffusion models (such as our work) - a current gap in the deepfake literature.
A.2 More results on cross-identity reenactment
Figure 6: Additional qualitative comparison with SOTA methods on cross-identity reenactment.
A.3 More results on out-of-domain reenactment
Figure 7: Additional qualitative comparison with SOTA methods on out-of-domain reenactment.
A.4 Limitations
Although our method provides high-fidelity results with excellent image quality, there are still some limitations. First, noticeable flickering occurs when generating videos. Second, the inadvertent generation of hands affects the quality of the generated content. Future studies can focus on addressing these issues.

### v2

@@ 4.4 Ablation Study
Table 3: Quantitative comparison with or without Facial Alignment and the Expression Adapter. The addition of both components allows for a better trade-off between identity characteristics consistency vs expression fidelity. We find that combining both provides the best qualitative and quantitative results.
. Same-identity Reconstruction Cross-identity Animation FA EA CSIM APD AED CSIM APD AED - - 0.75 0.011 0.098 0.52 0.021 0.182 - - - - 0.59 0.022 0.206 - 0.75 0.011 0.091 0.51 0.021 0.158 - - - 0.57 0.022 0.174
Figure 5: Qualitative comparison of results with or without Facial Alignment and the Expression Adapter. FA improves the generation quality, while EA maintains expression consistency. Ensembling both provides the best overall consistency. These observations are consistent with the quantitative results in Table 2.
We conduct an ablation study on the proposed 1) Facial Alignment (FA) and 2) Expression Adapter (EA). Facial Alignment is used exclusively for cross-identity animation. The baseline method, without FA and EA, is similar to the structure of AnimateAnyone, incorporating 2D surface normal snapshots from DECA. As illustrated in Table 3, adding FA significantly improves identity characteristics preservation (CSIM). FA uses the shape embedding from the reference image to render the 2D surface normal snapshots, thereby enhancing appearance preservation. Without it, the output tends to adopt the shape of the target motion rather than preserving the reference image’s features. This is particularly problematic when the reference and target images are from different styles, as it exacerbates the appearance mismatch, as shown in Fig. 5. On the other hand, FA ensures that the spatial scale of the surface normal snapshots matches that of the ReferenceNet, enhancing the spatial consistency of the generated output, especially when there is a substantial facial scale difference between the reference image and target motion. EA significantly enhances expression accuracy (AED) in both same-identity reconstruction and cross-identity animation. Methods incorporating EA can provide more expression-related details, and achieve greater accuracy in certain areas, such as the mouth and eyes, as depicted in Fig. 5. Combining both FA and EA can effectively handle the trade-off between appearance consistency and expression accuracy.
@@ 5 Ethical Considerations
We are committed to the positive applications of virtual stylized avatar animation in fields such as film, video conferencing, virtual education, and telemedicine. For instance, it can be used to safeguard privacy [65], protecting real identities in various digital environments, including social media, live streaming, telemedicine, and sensitive interviews. Acknowledging the broad applicability of these technologies, we are also aware of their potential for misuse when combined with other methods on photorealistic heads, as they cannot independently create complete videos with audio. We firmly oppose the creation of misleading or harmful content and aim to help promote the ethical use of these technologies. We recommend that all applications of these technologies incorporate watermarks into the images they produce. Furthermore, detection for generative content provides a practical solution to this issue, serving as an effective defender by identifying whether a video is real or generated. We are dedicated to contributing to this field by supporting the development of detectors. To this end, we evaluate the effectiveness of current state-of-the-art detectors and analyze their generalization capabilities, helping to identify potential areas for improvement and suggesting possible solutions in 5.1. This will serve as a foundation for future research aimed at mitigating weaknesses of existing detectors and providing benchmarking data for future approaches. Additionally, we are committed to strictly limiting our model’s use to ethical research purposes and actively supporting the detection community by sharing insights and resources. Privacy considerations can be found in 5.2. As responsible researchers, we call on other scholars to prioritize ethical considerations in their work. Additionally, our current method cannot produce videos with photorealistic avatars that resemble real-life footage without visible artifacts and flickering issues.
5.1 Generalization Analysis in Synthetic Face Detection
Stylized avatar animation is an important area of active research with numerous ethical applications. However, as with any generative technology, there are inherent risks related to potential misuse. Detection for synthetic face serves as a important safeguard by identifying generated content and distinguishing it from authentic footage, thus helping prevent the spread of misinformation and protecting individuals from potential harm. To address this risk, we conduct a detailed and comprehensive analysis of state-of-the-art detection techniques on our method and report their zero-shot performance results. This contributes to the understanding of the generalization capabilities of existing detectors, which is essential for advancing responsible applications.
Table 4: Detection frame-level and video-level AUC from the state-of-the-art detectors with respective pretrained weights on videos generated by our method AniFaceDiff. The best results are shown in bold.
@@ 5.1 Generalization Analysis in Synthetic Face Detection
We emphasize the importance of using generative technologies responsibly, ensuring that generated content can be effectively detected. To this end, we evaluated the videos produced by our method using pretrained, state-of-the-art detection techniques. Table 4 presents both frame and video-level detection results. Although video-level detection is more common, we include frame-level detection due to the possibility of temporally partial synthesized videos [74] i.e. having both real and generated segments in the same video. Our evaluation encompasses multiple types of detectors including naive detectors [66, 68], a frequency-level detector [70] and spatial detectors [67, 69, 71, 72, 73]. We notice that videos generated with AniFaceDiff had high detection rate by the tested detectors. Frame-level detection is more challenging since the detection performance depends on the prediction of each frame, rather than a majority voting across the frames in a video. Among the recent detection methods tested, CORE and CADDM achieve higher AUC scores, likely due to their enhanced feature learning capabilities. These results align with the specific features each method targets. CORE employs explicit regularization to ensure representation consistency across different augmentations, effectively mitigating overfitting. On the other hand, CADDM addresses implicit identity leakage, thereby improving its generalization on unseen data. Overall, enhancing the generalization capabilities of detection algorithms is crucial to mitigating overfitting and ensuring their effectiveness across different generative methods. Different characteristics from generative methods sometimes make it challenging for a single detector to perform optimally across all scenarios. By utilizing multiple detectors that employ different feature-learning strategies, we can better address the challenges posed by synthetic content and improve the reliability of detection, thus providing more comprehensive and robust safeguards against potential misuse.
@@
This assessment aims to pave the way for future studies on detection. Additionally, we aim to support the creation of higher-quality datasets for detection, particularly those generated using state-of-the-art diffusion models-based methods, addressing a current gap in the literature.
5.2 Privacy Considerations
Dataset Usage
This research utilizes the VoxCeleb and VoxCeleb2 datasets developed by the Visual Geometry Group (VGG) at the Department of Engineering Science, Oxford University, which come with an explicit privacy notice [^2] https://www.robots.ox.ac.uk/~vgg/terms/url-lists-privacy-notice.html. We use these datasets to evaluate the model, conduct benchmarking comparisons with previous methods, and assess detectors. Notably, we do not use any metadata, such as identity or nationality, in this research.
Protection of Privacy and Data Integrity
In our research, we recognize the importance of ethical considerations and privacy concerns. Our focus is on animating stylized avatars, which can be utilized to protect privacy in online environments. To address privacy concerns, we take several measures: First, all characters presented in this paper are entirely virtual and do not correspond to real individuals. Second, this study does not aim to identify or analyze personal data within publicly available datasets in any way that could lead to re-identification. We minimize the personal data by removing personal identity information (metadata) and voice data from the dataset, further preventing cross-correlation between data sources. Third, all images in the datasets are resized to 256x256, effectively reducing facial details and enhancing de-identification. Finally, access to the data, models, and model-generated outputs are strictly controlled. No private or personally identifiable information from individuals is collected or used in this research. All data processing adheres to privacy standards and does not involve identity recognition. We advocate for the responsible use of these technologies, emphasizing that they are applied in ways that uphold individual rights and contribute positively to society.
6 Conclusion
In this work, we introduce Anifacediff, a novel framework built on the Stable Diffusion model specifically designed for stylized avatar animation. The objective is to accurately maintain reference image consistency while effectively animate it with the pose and expression derived from the target motion video. To achieve this, we developed a new conditioning module that integrates two key components: Facial Alignment (FA) and Expression Adapter (EA). Facial Alignment ensures that the identity characteristics of the reference image are preserved and that spatial information is consistently aligned, while the Expression Adapter captures and conveys the intricate emotional nuances present in the target motion. Extensive experiments demonstrate that our method outperforms state-of-the-art techniques, especially when handling out-of-domain data across a wide range of styles. Additionally, our analysis of current state-of-the-art detectors evaluates their generalization capabilities, highlighting potential areas for improvement and possible solutions. This contribution reinforces our commitment to advancing detection techniques for generative content and promoting responsible development.
6.1 Limitations
Although our method delivers high-quality results, it has several limitations. First, flickering occurs during video generation. Second, the inadvertent generation of hands negatively impacts the quality of the produced content. Furthermore, the model’s performance is compromised if the character to be animated is overly abstract. Future studies could focus on addressing these issues. Finally, the low inference speed of diffusion models hinders the application in real-time scenarios. This limitation could be mitigated by developing faster samplers for diffusion models while maintaining high-quality generation.
Figure 6: Additional stylized animation results produced by our method.